OpenStudy (anonymous):

Laura fills a bucket with water at a rate of 6 L/min. At the same time, Sarah empties a bucket holding 30 L of water at a rate of 9 L/min. Let l represent the number of liters of water and let t represent time in minutes. The system models this situation. How many minutes will it take for the buckets to hold equal amounts of water, and how much water will that be? l = 6t l = 30 – 9t It will take ______ minutes for both buckets to hold equal amounts of water. They will each hold ______ liters.

OpenStudy (danjs):

V water = Flow rate * time is the unit in this prob

OpenStudy (danjs):

set them the same, and maybe there is a time value where that happens

OpenStudy (danjs):

one starts at 0 volume and Increases over time, the other starts at 30 and dereases over time , there will be some time when they both have the same volume of water

OpenStudy (anonymous):

dude it will be fine if u say the anwser

OpenStudy (anonymous):

dont want to take up your day

OpenStudy (danjs):

l = 6*t l = 30 – 9*t are the right relationships for the volume water in each bucket for any time t set them equal to see when they are the same, 6*t = 30 - 9t

OpenStudy (anonymous):

?

OpenStudy (anonymous):

...

OpenStudy (danjs):

one bucket starts empty and increases by 6 per min the other starts at 30 and looses 9 per min the linear equations for those are the ones they gave you , l = 6*t l = 30 – 9*t you set the amount of water in each bucket equal so you can say 6*t = 30 - 9*t solving that, t = 2 min

OpenStudy (danjs):

when the time is 2 min from the start, both have an equal amount of water, I = 6*2 = 12 Liters I = 30 - 9*2 = 12 Liters both have 12 Liters water at exactly the two min mark

OpenStudy (anonymous):

is that the anwser?

OpenStudy (danjs):

It will take ___two___ minutes for both buckets to hold equal amounts of water. They will each hold __twelve____ liters.

OpenStudy (anonymous):

ok then